OpenStudy (anonymous):

The church is 50 feet north and 100 feet east of the town center and the school is 100 feet north and 200 feet east of the town center. The coffee shop is the midpoint of these two places. How far north and how far east is the coffee shop from the town center?

OpenStudy (anonymous):

@gypsy1274

OpenStudy (anonymous):

pls help

OpenStudy (anonymous):

Plot these points on a graph. North/South will be the y coordinate and East/West will be the x coordinate.

OpenStudy (anonymous):

they dont hae south and west?

OpenStudy (anonymous):

North would be +y South (If that direction were given) would be -y East would be +x and West (if that direction were given) would be -x The fact that there are no south and west directions means it is easier because you are only dealing with positive numbers in the first quadrant. Draw it out and it will be easier to see.

OpenStudy (anonymous):

You will use the midpoint formula to find the location of the coffee shop. Then plot that point to check your answer.

OpenStudy (anonymous):

50,0 and 100,0

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Here is a good place to find the midpoint formula: http://www.purplemath.com/modules/midpoint.htm

OpenStudy (anonymous):

Your answer should be one ordered pair (x,y)

OpenStudy (anonymous):

\[\frac{x_1 + x_2}{2},\frac{y_1+y_2}{2}\]

OpenStudy (anonymous):

x1=50 + x2=100/2 and y1=100 y2=200/2 ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

75,150? is that the answer

OpenStudy (anonymous):

Yes, that is the answer I got as well.

OpenStudy (anonymous):

ok thank you soo much yur alot of help

OpenStudy (anonymous):

Your welcome.